Prev Next

SW&IT Meeting Minutes 2021-11-22

Attendees: Brandon, David, Glenn, Matthieu, Paul, Stephen, Xianyu, Yang


Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

LR20211115 - Bec, Matthieu 40.0
LR20211115 - Mechtley, Brandon 80.0
LR20211115 - Eychaner, Glenn 40.0
LR20211115 - Edmeades, David 40.0
LR20211115 - Grenz, Paul 12.0
LR20211115 - Hooper, Stephen 0.0
LR20211115 - Zhang, Yang 40.0
LR20211115 - Zhang, Xianyu 40.0

News from management

Previous Action Items

Owner Action Due Date Completion Date

Open topics

Team reports


  • Last week
    • OC, Management
    • safety+cybersecurity trainings
    • learning some more about AO s/w, join discussions and scheming on #ao-sw
    • LBTplot conversion to DMS (ctd.)
    • PB s/w discussions (ctd)
    • logrotate mongodb
    • ao-supervisor and git/github auth on shared machine
    • shark-nir PAE
  • This week
    • Performance Appraisals
    • git/github "policy" (= philosophy slides) for new staff
    • shark-nir PAE ctd.


  • Previous:
    • Met with Olga and discussed several things:
      • We sorted out the problem with "IRAF/lbcrangebal" on obs5 (so it is good now).
      • We discussed MODS WCS calculation and agreed on how to compute various things (will get together later to do more experiments).
      • We also discussed some TMS operator experiences and troubles (will have to think more on that).
    • Some TCS/AGW1 troubleshooting on Tue (Josh called on Friday evening for AGW1/IIF problems, I was able to resolve it for him before opening).
    • AOS Patch activated (finally) and no problems reported.
    • TMS meeting and some discussions:
      • Andrew plans on reviving the missing channel solution (will need to work with him on this).
      • Some utilities and enhancements may be needed.
      • Automatic oordination between TMS and other software (e.g., FPIA) seems difficult at this time (may need new infrastructure support).
  • Next:
    • Vacation + Holidays (Happy Thanksgiving to everyone!)


  • Last week
    • Had Stomach Flu, was out of commission for two days.
    • Finishing up proxy code and testing.
    • Proxy works, still need to test with both processes talking through it.
  • This week
    • Finish testing proxy.
    • Thanksgiving!


  • Last week
    • Progress on Win Server, but I have trust issues regarding license keys and Autodesk rules
    • Synology equipment looks good, off-label drives get a whine but they work. I am anticipating that at some point they will be on the approved list. I will head in tomorrow and get the units set up to the point that the drives can burn in and we can remotely configure the software
    • UITS bug(?) with core switching hit my office, deactivating my Ethernet connection. My contact at NetOps was able to fix it with only the room number, so if you try to remote into a system and find that it's unresponsive let me know and I can have him check it out.
    • Gear ordered for AB, long lead times for everything.
    • VMWare licenses ordered
    • Discussion, conclusion on new PDUs for Tucson
    • Moved Luci WebIO units + power adapter + Dell power supply into GE's office
    • Ticket faux pas re: mask cutter UPS battery replacement, sorry for making it complicated.
    • Jim Word issue, not entirely sure what's up with that
    • XZ data move, warnings about space.
  • This week
    • Win Server exploration tree (New 2016 serial, see if Vault 2020 will work on Server 2019 [preferred]), etc.
    • Synology build/burn in
      • Fun fact: we are getting close to "losing" 2TiB on drives when you consider their decimal nameplate statement vs. the binary capacity. 1TB * .909495 = TiB, so a 22TB nameplate drive is ~20TiB usable capacity. 20TB drives are available now.
    • See if I can remote into Jim's system and see what is up with Word.
    • Per MS request, spec new laptop for Jim
    • Sundries


  • Last week
    • Digging through the 550,000 lines of code of the LUCI system.
    • Created a new account on LUCI for Jared Schuh for mask exchanges and other instrument work.
    • Tested the new LUCI software live during engineering. Worked out how to save SRR mode images with Dave Thompson. Performed more tests in simulation.
  • This week
    • Digging through the 550,000 lines of code of the LUCI system.
    • Create a release version of the new LUCI software and place it in stow on the observer machines.
    • Continue working on the RTD issues.
    • Examine the other LUCI GUIs to check for GUI modifications from threads other than the Event Dispatch Thread.


  • Last week
    • PB upgrade:
      • Met with Marco
      • Made plan/progress spreadsheet, annotating PB_SW_Plan.docx
    • SOUL commissioning meetings
    • Deployment:
      • ao-supervisor new stable/test directories ready
      • Made {use,test}_{uao,soul} scripts.
      • Put a draft of a development guide on the GH wiki, based on conversation with MB, XZ.
    • hk-partial-timeout ready to test, made testing scripts
    • Testing SOUL-devel (latest) with Juan Carlos
      • Merged new look up tables from Enrico (need to get these in GH)
      • Pupil rerotator problems
      • Acquisition camera viewer closed immediately on opening, need to test more
      • FLAOGUI doesn't show plots if there are skip frames (plots from steps before skip frame don't seem to be produced?)
      • Problems with MsgD routing (@M_ADSEC syntax) not working, needed to restart msgdrtdb process
  • This week
    • AO data migration
    • Testing with Juan Carlos (SOUL-devel latest; hk-partial-timeout and config-gain-ramp if time)
    • LBTI issues if time
    • Vacation day next Monday


  • VM discussion
  • Still need quotes for Cisco
  • OAC stuff
  • PLC stuff briefly
  • Workshop Wednesday, Thursday


  • VMware install on mountain and Tucson EPYC
Topic revision: r10 - 22 Nov 2021, DavidEdmeades
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback